OpenStudy (anonymous):

Please help! For the equation x^2 + 3x + j = 0, find all the values of j such that the equation has two real number solutions.

OpenStudy (turingtest):

what is the formula for the discriminant, which will tell you the properties of the roots?

OpenStudy (anonymous):

wouldn't it be 4x + j = 0?

OpenStudy (turingtest):

not quite in the quadratic form\[ax^2+bx+c=0\] the discriminant is given by\[b^2-4ac\]what is that in your case?

OpenStudy (anonymous):

I'm still lost

OpenStudy (turingtest):

you have x^2 + 3x + j = 0 the general form is ax^2+bx+c=0 so what are the coefficients a, b, and c in your case?

OpenStudy (turingtest):

compare\[ax^2+bx+c=0\]\[x^2+3x+j=0\]what is \(a\) in the second formula? what is \(b\) ? what is \(c\) ?

OpenStudy (anonymous):

a is x, b is 3 and c is j

OpenStudy (turingtest):

almost; b and c are right, but... a is the coefficient of x^2 if x^2 is by itself, what is its coefficient?

OpenStudy (anonymous):

x

OpenStudy (turingtest):

remember that\[x^2=1x^2\]so\[a=1\]

OpenStudy (turingtest):

so now you can use the formula for the discriminant I wrote above what will it be in your case?

OpenStudy (anonymous):

so it's 1x^2 + 3x + j = 0?

OpenStudy (turingtest):

well, that's what you are given with the 1 written out explicitly (which is not really necessary) and what is the formula for the discriminant that I wrote above?